### v3

# Enhancing Transformers for Generalizable First-Order Logical Entailment

Abstract
Transformers, as the fundamental deep learning architecture, have demonstrated great capability in reasoning. This paper studies the generalizable first-order logical reasoning ability of transformers with their parameterized knowledge and how to improve it. Transformers’ capability of first-order reasoning is further captured by whether they can conduct first-order logical entailment, which is quantitatively measured by their performance in answering knowledge graph queries. We establish the connections between (1) two types of distribution shifts studied in out-of-distribution generalization and (2) unseen knowledge and query settings discussed in the task of knowledge graph query answering, which makes it possible to characterize the fine-grained generalizability. Results on our comprehensive dataset showed that transformers outperform previous methods designed particularly for this task and provided detailed empirical evidence about the impact of the input query syntax, token embedding, and transformer architectures on their reasoning capability. Interestingly, our results revealed the mismatch of positional encoding and other design choices of transformer architectures in previous practices. Motivated by this, we propose TEGA, a logic-aware architecture that significantly improves the performance in generalizable first-order logical entailment.
[…]
Enhancing Transformers for Generalizable First-Order Logical Entailment
[…]
1 Introduction
As a fundamental architecture in deep learning, transformers possess strong reasoning capabilities on various tasks, including arithmetic reasoning Saxton et al. (2019); Hendrycks et al. (2021), symbolic reasoning for first-order logic rules Dehghani et al. (2019); Lample and Charton (2019), set-theoretic operations Barrett et al. (2018), and theorem proving Polu and Sutskever (2020). Besides, transformers have also demonstrated proficiency in logical reasoning over natural language Han et al. (2022); Tian et al. (2021). To distinguish whether transformers conduct the reasoning rather than fitting the data distribution, recent studies in natural language reasoning further measure the capabilities of transformers for out-of-demonstration samples (Saparov et al., 2023). However, their discussions are limited in two aspects: (1) they only concern the reasoning ability of transformers with in-context knowledge, and (2) they fail to elicit the connection of out-of-demonstration samples with two distribution shifts (Moreno-Torres et al., 2012) for the study of out-of-distribution generalization.
In this paper, we take a step further to understand the generalizable reasoning capability of transformers. What sets us apart from (Saparov et al., 2023) is that (1) our investigation covers the transformer reasoning with parameterized knowledge, which suits many scenarios when the related knowledge is not explicit for users and only questions are given, and (2) we further realize two distribution shifts (Moreno-Torres et al., 2012) in our first order reasoning tasks, which boils down to the process of first-order logical entailment, where we verify whether a first-order sentence (conclusion) is the logical consequence of a set of known first-order sentences (premises) or not (Marker, 2002).
Specifically, we study the first-order logical entailment with Knowledge Graphs (KG), leading to the widely discussed task of knowledge graph query answering. In our setting, knowledge in KGs is parameterized in models and forms implicit premises. (§2) The process of logical entailment occurs when identifying the answer to a logical query. Therefore, the two distribution shifts – concept shift and covariant shift – are naturally credited to the unobserved knowledge and the unseen query types in our experimental settings. (§3) For better evaluation, we build our own benchmark with fifty-five types of logical queries over three knowledge graphs (FB15k Bordes et al. (2013), FB15k-237 Toutanova and Chen (2015), and NELL995 Carlson et al. (2010)), including all existing features discussed in recent literature (Yin et al., 2023b). Our benchmark results suggest that transformers can handle first-order logical entailment even compared to the methods particularly designed for this task. (§4)
We conducted extensive experiments to characterize the impact of three main aspects of solving logical entailment on reasoning capability: the query syntax, the learning of token embedding, and the transformer architecture. Our results provide strong empirical clues on improving transformers’ reasoning capability by choosing the proper formal language syntax, positional encodings, semantics in pre-trained KG embeddings, and inductive biases in transformer architectures. (§5)
Interestingly, our results demonstrated the superior performance and generalizability of relative positional encoding (RPE) over traditional absolute positional encoding (APE). However, previous studies have only proposed inductive biases within the APE setting, which we prove ineffective in the RPE setting. To fill this gap, we propose TEGA (Transformer Encoder with Guided Attention), a novel modeling methodology that facilitates effective reasoning with logic-aware guidance in self-attention. Our study shows that TEGA substantially improves the performance and generalizability upon transformers under the RPE setting. (§6)
Overall, this work makes three key technical contributions:
  1. 1.
    Benchmark. We conducted comprehensive benchmarking of transformer architectures against baseline models for knowledge graph query answering, using a new dataset that incorporates two distribution shifts.
  2. 2.
    Exploration. We systematically investigated critical design choices throughout the modeling process, including query syntax and permutation, token embedding, and transformer architecture designs.
  3. 3.
    Method. We introduced TEGA, a novel architecture that improves logical entailment performance and generalizability through two inductive biases. We validated its effectiveness through extensive experiments and ablations.
Our code and data can be found at https://github.com/HKUST-KnowComp/TEGA.
2 Preliminaries
This section briefly introduces First-Order (FO) logical entailment and its restricted application in knowledge graph query answering. By revealing the connection, we show that the key ability to address complex queries is the first-order logical entailment. For simplicity, our introduction only restricts to a finite entity set $\mathcal{E}$ and a finite binary relation set $\mathcal{R}$. No functions are mentioned. Detailed presentation of first-order logic can be found in model theory literature (Marker, 2002).
2.1 First-order Logical Entailment
Definition 1 (First-order sentence).
The set of first-order formulae is the minimal set $F$ such that
  1. 1.
    $p(s,o)\in F$, where $p$ is the relation, $s,o$ can be either an entity in $\mathcal{E}$ or a variable.
  2. 2.
    If $s\in F$, then $\lnot s\in F$; If $s,t\in F$, then $s\land t\in F$ and $s\lor t\in F$.
  3. 3.
    If $s(x)\in F$ and $x$ is a variable that is not qua- ntified, then $\forall x.s(x)\in F$ and $\exists x.s(x)\in F$.
We say a variable $x$ in a formula is quantified if there is a quantifier ($\exists x$ or $\forall x$). Otherwise, we say a variable is free. A sentence is a formula without free variables.
Next, we introduce first-order logical entailment, the central concept of first-order reasoning. In general, the entailment is a process of verifying the conlusion (one sentence to verify) given the knowledge (a set of given premises). Notably, knowledge, conclusion and verification are the three key components of the definition. For first-order logical entailment, the knowledge and conclusion are restricted as first-order sentences, and the verification process is subject to the first-order logical calculus. A set of FO sentences $\{p_{i},i=1,...,n\}$ entails a FO conclusion $s$ is denoted as:
$\underbrace{\{p_{i},i=1,...,n\}}_{\text{ {knowledge} or premises}}\models\underbrace{s}_{\text{{conclusion}}}$ (1)
After verification, its truth value suggests whether $s$ is the logical consequence of the premises.
2.2 Knowledge Graph Query Answering
Knowledge graph query answering is an important application of FO logical entailment. Its importance comes from both the importance of knowledge graphs and the queries in database systems. Here, we connect the task of KG query answering and FO entailment from the aspects of knowledge, conclusion, and verification.
Knowledge. Given an entity set $\mathcal{E}$ and relation set $\mathcal{R}$, a knowledge graph is defined as a set of triplets $\mathcal{G}=\{(h_{i},r_{i},t_{i})\}$, where $h_{i},t_{i}\in\mathcal{E}$ are entities and $r_{i}\in\mathcal{R}$ are relations. Each triple $(h,r,t)$ in KG can be regarded as the simplest form of first-order sentence $r(h,t)$. Thus, the knowledge graph $\mathcal{G}$ provided the premises.
Answers to the query as multiple conclusions to verify. Existing research for KG query answering usually focuses on the restricted families of FO queries, specifically, the Existential First-Order (EFO) queries. An EFO query is represented in Disjunctive Normal Form (DNF) as a disjunction of one or multiple conjunctions:
$q[V_{?}]=V_{?}.\exists V_{1},V_{2},...,V_{n}:c_{1}\vee c_{2}\vee...\vee c_{m}$ (2)
In equation 2, $V_{?}$ is the free variable, $V_{1},V_{2},...,V_{n}$ are existential variables, while $c_{1},c_{2},...,c_{m}$ are conjunctions of one-hop atomic formulas or their negation, defined as $c_{i}=\alpha_{i1}\wedge\alpha_{i2}\wedge...\wedge\alpha_{ik_{i}}$ and $\alpha_{ij}=r(h,t)$ or $\neg r(h,t)$, where $r\in\mathcal{R}$, $h$ and $t$ are either constant entities $e\in\mathcal{E}$ or variables.
We can see that an EFO query is not a FO sentence because it contains a free variable to query. The goal of answering a query is to find all possible entities $a$ such that $\mathcal{G}\models q[a/V_{?}]$, where $q[a/V_{?}]$ is a FO sentence where the free variable $V_{?}$ is substituted by entity $a$. To summarize, the answer set
$A(\mathcal{G},q)=\{a\in\mathcal{E}:\mathcal{G}\models q[a/V_{?}]={\rm True}\}.$ (3)
Next, we show that the key condition to be verified, i.e., $\mathcal{G}\models q[a/V_{?}]={\rm True}$, relies on the knowledge graph $\mathcal{G}$. By FO logical calculus, evaluating $\mathcal{G}\models q[a/V_{?}]$ is equivalent to:
$\bigvee_{V_{1},...,V_{n}\in\mathcal{E}}\left[\bigvee_{i=1}^{m}\left[\left(\mathcal{G}\!\models\!\alpha_{i1}\right)_{V_{?}=a}\wedge...\wedge\left(\mathcal{G}\!\models\!\alpha_{ik_{i}}\right)_{V_{?}=a}\right]\right].$ (4)
We could see that whether $\mathcal{G}\models q[a/V_{?}]$ is True essentially depends on many times of FO logical entailment on atomic formulae $\mathcal{G}\models\alpha_{ij}$, where each of these depends on whether the atomic $r(h,t)$ associated to $\alpha_{ij}$ is contained in $\mathcal{G}$ (or not in $\mathcal{G}$ for the negation case).
Figure 1: An illustration of our two-fold Out-of-distribution Generalization in knowledge and query type dimensions. Figure (a) explains the relation between OOD shifts in the two dimensions. Figure (b) and Figure (c) provide examples demonstrating the causes of OOD shifts in the two dimensions.
3 Distribution Shifts in KGQA
This section introduces the formulation of distribution shifts in KG query answering tasks. Such tasks are connected to first-order reasoning due to their close connection with logical entailment. By clearly presenting their connection with distribution shifts studied in statistical ML community (Moreno-Torres et al., 2012), they are then able to measure the generalizable first-order reasoning capability of deep learning models, in particular transformers, with parameterized knowledge.
Problem formulation. A dataset for KG query answering is denoted as a set of samples $\{(x_{i},y_{i})\}$ drawn from the joint distribution $P(X,Y)$. Here, $X$ represents the random variable corresponding to an EFO query, with $x_{i}$ as individual samples, and $Y$ represents the random variable for the answer set, with $y_{i}$ as individual samples.
This paper studies transformers, where the input $X$ is tokenized and fed into the transformer architecture, and the output $Y$ is predicted with a classifier, where all entities in the KG are ranked by embedding similarities. In both the training and testing phases, the knowledge graph $\mathcal{G}$ is not explicitly accessed by the model. In this way, the necessary condition for transformers to correctly answer the queries is that the knowledge in KG is materialized into their parameters. To examine whether transformers possess the generalizable first-order reasoning capability, we consider two classic types of out-of-distribution shifts: concept shift and covariant shift. They are connected to unobserved knowledge and unseen query types.
3.1 Two Types of Distribution Shifts
Out-Of-Distribution (OOD) generalization is one of the key topics in machine learning research Vapnik (1991); Quionero-Candela et al. (2009). It considers the shifts in the joint distribution $P(X,Y)$ from the training to the testing phases. Equation 5 decomposed the distribution shift by applying conditional probability formula, introducing two basic types of OOD shifts: concept shift refers to the changes in $P(Y|X)$, while covariate shift refers to the changes in $P(X)$.
$P_{\rm train}(Y|X)$ $P_{\rm train}(X)\neq\underbrace{P_{\rm test}(Y|X)}_{\rm Concept}\underbrace{P_{\rm test}(X)}_{\rm Covariate}$ (5)
As illustrated in Fig. 1, we studied distribution shifts in both knowledge and query type dimensions.
Concept Shift by Unobserved Knowledge
In our KG query answering setting, we construct an observed knowledge graph $\mathcal{G}_{o}$ as a subset of the full knowledge graph $\mathcal{G}$. In the training stage, the models only access the answers from the observed graph, denoted as $A(\mathcal{G}_{o},q)$. While in the testing stage, they are evaluated on answers from the full graph, denoted as $A(\mathcal{G},q)$. This setup introduces a concept shift in the conditional distribution $P(Y|X)$ between training and testing because the change of knowledge causes different results of entailment, see Equation (3). Therefore, the set difference $\mathcal{A}_{ood}=A(\mathcal{G},q)-A(\mathcal{G}_{o},q)$ contains the entities that can only be derived when the model generalizes from knowledge $\mathcal{G}_{o}$ to $\mathcal{G}$. By measuring the performances of models on $\mathcal{A}_{ood}$, we are able to access how a model generalizes under the concept shift caused by unobserved knowledge. This is also termed knowledge generalization or knowledge inference in the literature Sun et al. (2021).
Dataset Operations Features Unseen Types
pro. int. uni. neg. exi. mul. cyc.
GQE 0
Q2B 4
BetaE 4
FIT 10
SQE 29
Ours 32
Table 1: Comparison between different benchmarks on their supported logical operations, query features, and the number of unseen query types.
Covariate Shift by Unseen Query Types
We also examine the capability of query answering models to generalize and maintain performance on query types that are unseen in the training stage. In our testing data, we include new query types that are combinatorially crafted from the training query types, introducing a covariate shift in the distribution of $P(X)$. The models’ performances on these unseen types are measured to evaluate their generalizability to covariate shifts in the query type dimension. This is also termed combinatorial generalization or compositional generalization in the literature Wang et al. (2021); Yin et al. (2023a).
4 Evaluation Benchmark
To better investigate the performance in generalizable FO reasoning, we build a new benchmark of KG query answering for our experiment. We first introduce the dataset construction process and evaluation settings, and then present our results of transformers in comparison to previous methods.
Dataset Model Type ID (Q) OOD (Q) All Queries
ID (K) OOD (K) ID (K) OOD (K) ID (K) OOD (K)
FB15k BetaE Ren and Leskovec (2020) Probabilistic 26.9 18.5 22.4 13.5 24.3 15.6
ConE Zhang et al. (2021) Geometric 35.5 22.0 27.2 15.6 30.7 18.3
CQD Arakelyan et al. (2021) Fuzzy Logic 33.1 20.7 21.5 11.2 26.4 15.2
LMPNN Wang et al. (2023b) GNN 32.5 21.0 24.3 13.9 27.7 16.9
[…]
Trans.+Absolute PE Transformer 46.9 31.9 21.8 13.2 32.3 21.0
Trans.+Disentangled PE Transformer 51.7 33.3 23.7 13.6 35.4 21.8
[…]
FB15k-237 BetaE Ren and Leskovec (2020) Probabilistic 27.4 16.7 23.3 13.1 25.0 14.6
ConE Zhang et al. (2021) Geometric 34.9 18.5 27.3 14.2 30.5 16.0
CQD Arakelyan et al. (2021) Fuzzy Logic 32.8 14.1 20.2 8.6 25.5 10.9
LMPNN Wang et al. (2023b) GNN 26.6 16.4 22.3 12.4 24.1 14.1
[…]
Trans.+Disentangled PE Transformer 59.0 20.0 21.3 8.7 37.0 13.4
[…]
NELL995 BetaE Ren and Leskovec (2020) Probabilistic 40.9 16.7 32.3 11.1 35.9 13.4
[…]
CQD Arakelyan et al. (2021) Fuzzy Logic 42.9 14.7 28.9 8.3 34.8 11.0
LMPNN Wang et al. (2023b) GNN 42.8 17.1 32.1 11.6 36.6 13.9
[…]
Trans.+Absolute PE Transformer 68.8 18.9 25.8 8.3 43.8 12.7
Trans.+Disentangled PE Transformer 73.6 19.3 29.7 8.7 48.1 13.1
[…]
Table 2: Experimental result (in MRR%) of Transformers in comparison with previous state-of-the-art methods with different backbones in our full benchmark. Definitions of evaluation metrics are provided in Sec. 4.2.
4.1 Dataset Construction
Table 1 compares the feature and statistics of query types selected in our benchmark with previous benchmarks in KG query answering, namely GQE dataset Hamilton et al. (2019), Q2B dataset Ren et al. (2020), BetaE dataset Ren and Leskovec (2020), FIT dataset Yin et al. (2023b), and SQE dataset Bai et al. (2023b). We included the entire set of set operations and query features in our selection of 23 seen query types. Moreover, we compositionally crafted 32 unseen query types that are only used in the testing stage to examine models’ OOD generalizability in the query type dimension. For instance [^1] Queries are represented in EFO syntax Yin et al. (2023b)., suppose a query answering model is trained on query type 2in [(r1(s1,f))&(!(r2(s2,f)))] and 2p [(r1(s1,e1))&(r2(e1,f))], we expect it can also handle the unseen query type in2p [(r1(s1,e1))&(!(r2(s2,e1)))&(r3(e1,e2))&
(r4(e2,f))
], which is compositionally designed based on the previous two. All queries are sampled from the following three knowledge graphs: FB15k Bollacker et al. (2008); Bordes et al. (2013), FB15k-237 Toutanova and Chen (2015), and NELL995 Carlson et al. (2010). The details of all query types and their statistics in our benchmark are provided in Appendix A. We also provide query graphs for graph-augmented methods Liu et al. (2022); Xu et al. (2023) in our experiments, with detailed definitions in Appendix F.
4.2 Evaluation Settings
In accordance with the two types of distribution shifts discussed above, we evaluate the performance of query answering models across the two corresponding dimensions using the mean reciprocal rank (MRR) metric. In the knowledge dimension, we employ the notation ID (K) and OOD (K) to represent the performance on the answer set $\mathcal{A}_{id}=A(\mathcal{G}_{o},q)$ and $\mathcal{A}_{ood}=A(\mathcal{G},q)-A(\mathcal{G}_{o},q)$, respectively. Concurrently, in the query type dimension, we utilize the notation ID (Q) and OOD (Q) to represent the performance on seen query types and unseen query types, respectively. Higher scores in ID (K) and ID (Q) reflect the models’ effectiveness in performing logical entailment on in-distribution data, while higher scores in OOD (K) and OOD (Q) demonstrated better generalizability of the models. It is also noteworthy to mention that settings in ID (K) and OOD (K) denote the concept shifts, while those in ID (Q) and OOD (Q) capture the covariate shift. For details regarding the metric calculation, please refer to Appendix G.
Figure 2: An illustration of three core stages in the pipeline of KG query answering with transformers.
4.3 General Benchmarking Results
We evaluate transformer encoders [^2] Decoder-only and Encoder-Decoder yield comparable performance as illustrated in Appendix H. against five existing methods for KG query answering, with details provided in Appendix B. Meanwhile, we compare four positional encoding settings in transformers: Absolute PE is the default sinusoidal positional encoding introduced in the original transformer Vaswani et al. (2023). Relative PE Shaw et al. (2018) applied a learnable matrix based on the relative position of tokens. Disentangled PE He et al. (2021) uses separate vectors to encode the content and position of each token, decoupling the contextual and positional information during self-attention. Rotary PE Su et al. (2023) encodes position information by rotating query/key vectors in the complex space, capturing relative positions through rotation angles.
The experimental results are presented in Table 2. Transformers marginally outperform all baselines in both ID (K) and OOD (K) in seen query types. Both APE and DPE transformers have poor OOD generalizability in the query type dimension. In contrast, RPE and RoPE transformers demonstrates much stronger generalizability, as it can better capture structural information from the distances between tokens and flexibly handle query sequences of unseen lengths. Overall, RPE transformers achieve the best performance and OOD generalizability on first-order logical entailment.
5 Transformer Architectures and First-order Reasoning Capability
In this section, we further investigate the dependencies between transformers’ performance and multiple design choices throughout the modeling process. Figure 2 depicts the three-stage pipeline for deriving the answer set $\mathcal{A}$ (output $Y$) of query $q$ (input $X$) using transformers. We present the impact of query syntax in Section 5.1, token embeddings in Section 5.2, and transformer architecture in Section 5.3. We summarize the findings in Section 5.4.
5.1 Study on Query Syntax
Logical queries with the same semantics can be represented in different syntaxes with different formal languages. To investigate the impact of difference in formal languages, we conducted experiments on two formal languages: Lisp-like Syntax (Wang et al., 2021) represents query with fully parenthesized nested formula; EFO Syntax Yin et al. (2023b) represents query with all one-hop atomics connected with conjunctions and disjunctions in parallel. As the Lisp-like syntax has a limited representation scope, we used a subset of our benchmark that includes 13 seen query types and 12 unseen query types that are supported by both languages. The experimental results on the impact of formal language are presented in Table 3. In APE Transformers, there is no difference between the results of the two formal languages. However, with RPE, the EFO syntax achieves much better generalizability in the query type domain. This phenomenon can be explained by the differences in the structural features of the two languages: with a parallel structure, the distances (i.e., relative positions) between tokens with the same logical relationship are more consistent than those in a nested structure. As a result, RPE can better learn the logical relationship between tokens and better generalize to unseen query types.
Model Syntax ID (Q) OOD (Q)
ID (K) OOD (K) ID (K) OOD (K)
Trans.+Absolute PE Lisp-Like 57.1 20.7 10.0 4.9
EFO 58.1 20.7 10.4 5.1
Trans.+Relative PE Lisp-Like 58.3 20.9 22.1 9.8
EFO 58.1 20.9 35.4 14.3
Table 3: Experimental results of Transformers in FB15k-237 with different formal languages. Details of the dataset fragment are provided in Appendix A.
Model Before Reversion After Reversion
ID(K) OOD(K) ID(K) OOD(K)
Trans.+Absolute PE 54.1 18.3 27.8 9.8
Trans.+Relative PE 54.3 17.0 54.5 16.8
Table 4: Experiment results of Transformers on five query types before and after reversing permutation.
Furthermore, even for the same query in EFO Syntax, there can be various permutations of one-hop atomics, with the total number of potential permutations growing factorially with the length of the query. Therefore, it is crucial for transformers to robustly handle queries with different permutations. To evaluate this property, we selected five query types (2p, 3p, ip, pi, 2in) and reversed their permutations while maintaining the same semantics in the testing data (see Appendix C). The experimental results on query permutation are presented in Table 4. We observed a significant performance drop after reversing the query in APE transformers. However, for RPE transformers, the performance before and after reversing the permutation remains consistent, demonstrating their robustness in handling queries with modified permutations.
5.2 Study on Token Embeddings
Some methods in KG query answering Arakelyan et al. (2021); Wang et al. (2023b); Xu et al. (2023) utilized pre-trained KG embeddings from link predictors in their query encoding process, while other methods Ren and Leskovec (2020); Zhang et al. (2021); Bai et al. (2023b) randomly initialize the embedding matrix and jointly learn the KG embedding and model parameters from the training process. It’s still unclear that to what extent can pre-trained KG semantics can assist the performance of query answering models. To rigorously assess the impact of pre-trained semantics in token embeddings, we have selected three widely recognized pre-trained KG embeddings: TransE Bordes et al. (2013) models relationships by interpreting them as translations within the embedding space, DistMult Yang et al. (2015) employs a bilinear model wherein entities are represented as vectors and relations as diagonal matrices, ComplEx Trouillon et al. (2016) extends this framework into a complex vector space, allowing for more nuanced inter-entity interactions. In our experiment, all three KG embeddings are frozen during the training stage, in comparison to a baseline with randomly initialized and jointly learned embeddings.
The experimental results are presented in Table 5. Two stronger KGEs, ComplEx and DistMult, can effectively improve the performance of transformer. In contrast, the performance of TransE consistently lags behind the baseline across all three datasets. This discrepancy can be attributed to the baseline model’s embedding matrix learning process, which can be viewed as implicitly training a link predictor using a transformer architecture (e.g., KG-BERT Yao et al. (2019)). This modern approach may naturally outperform earlier models like TransE.
Dataset KGE ID (Q) OOD (Q)
ID (K) OOD (K) ID (K) OOD (K)
FB15k Random 48.1 32.3 35.4 21.5
TransE 44.3 30.3 34.2 20.8
DistMult 49.7 32.6 37.5 21.8
[…]
FB15k-237 Random 54.8 20.0 37.7 15.8
TransE 54.5 20.0 35.0 14.9
[…]
NELL995 Random 68.0 18.7 49.8 13.4
TransE 68.5 18.6 48.7 12.5
DistMult 72.2 19.9 51.2 12.9
[…]
Table 5: Experimental results of Transformer + Relative PE with different pre-trained KGEs.
5.3 Study on Transformer Architecture
Some existing literature Kotnis et al. (2021); Liu et al. (2022); Xu et al. (2023) has explored the use and design of transformers in KG query answering, in which several inductive biases in transformer architecture were proposed. However, as all these methods focused on different aspects and were evaluated on different benchmarks, a clearer investigation is needed. We implemented two designs from previous works on top of transformers: Adjacency Matrices Masking: an inductive bias proposed in kgTransformer Liu et al. (2022), which aims to enhance reasoning performance by limiting the range of self-attention to one-hop neighboring nodes in the query graph. Directed Distance Encoding: another inductive bias, introduced in Query2Triple Xu et al. (2023), that facilitates information aggregation by injecting a directed distance message into the self-attention mechanism. The experimental results on existing transformer designs are presented in Table 6. Both inductive biases effectively improve generalizability in the query type dimension when using APE. In contrast, under RPE, these improvements are not observed, and the use of adjacency matrices masking even results in a significant decrease in performance. This decrease is attributed to the masking in self-attention, which restricts message passing to only nodes within one hop, thereby limiting the model’s ability to perform complex logical reasoning. Regarding directed distance encoding, the information about directed distances is mostly contained in the relative positions of the linearized sequence, thus the improvements seen with original APE are not sustained.
Model ID (Q) OOD (Q)
ID (K) OOD (K) ID (K) OOD (K)
[…]
+Adjacency Matrices Masking 51.3 20.1 28.3 13.0
[…]
Table 6: Experimental result of Transformer + Absolute/Relative PE in FB15k-237 with existing inductive biases. Considering its triviality and compatibility with our dataset, the inductive bias from BiQE is not included in our experiment.
5.4 Summary of Empirical Findings
In terms of query syntax, we discovered that employing formal language with parallel structure can significantly enhance OOD generalizability in the query type dimension for RPE transformers. Moreover, we observed that transformers with relative PE can robustly handle queries with different permutations. Regarding transformer architecture, we found that the inductive biases proposed in previous approaches are only effective under APE, but fail to improve upon RPE transformers, which we have shown to be the preferred setting. This mismatch between transformer design and positional encoding settings motivated us to design a methodology that effectively boosts the performance and generalizability of RPE transformers in first-order logical entailment.
6 Transformer Encoder with Guided Attention
One of the main challenges in KG query answering is to perform multi-hop logical reasoning following the one-hop atomics and the logical operations between them. In traditional reasoning tasks, such as language-based reasoning, transformer models perform multi-hop reasoning by procedurally forming “reasoning trees” within their attention mechanisms Hou et al. (2023); Murty et al. (2023); Wang et al. (2023a). While in our task, it is challenging for transformers to capture the pattern of reasoning trees behind the queries, as the semantics of logical operators are neither explicitly modeled nor pre-trained, but implicitly learned from limited training queries. To navigate this conundrum, we present TEGA (Transformer Encoder with Guided Attention), a modeling methodology that boosts logical reasoning performance by providing logic-aware guidance within the self-attention mechanism of transformers. The following sections will introduce the two inductive biases in TEGA and their effectiveness with experimental results.
6.1 Logic-aware Relative Positional Encoding
Our analysis so far shows that RPE is the best positional encoding for transformers in our task for its superior performance and generalizability in both knowledge and query type dimensions. However, it still has limited expressiveness in the logical relationship between tokens, as the tokens with the same relative distance can have different logical relationships among queries. To fill this gap, we propose LogiRPE, a logic-aware mechanism to enhance self-attention. Specifically, for a sequence input sequence embeddings $x_{i}\in\mathbb{R}^{d}$, $i=1,...,n$, the enhanced self-attention is computed by
$z_{i}=\sum_{j=1}^{n}\alpha_{ij}(x_{j}W^{V}+\beta_{ij}^{V}),\quad\alpha_{ij}=\frac{\exp{e_{ij}}}{\sum^{n}_{k=1}\exp{e_{ik}}}$ (6)
$\quad e_{ij}=\frac{x_{i}W^{Q}(x_{j}W^{K}+\beta_{ij}^{K})^{T}}{\sqrt{d}}.$ (7)
where $W^{Q},W^{K},W^{V}\in\mathbb{R}^{d\times d}$ are the weight matrices for query, key, and value while $\beta_{ij}^{V},\beta_{ij}^{K}\in\mathbb{R}^{d}$ are two logical bias terms introduced by LogiRPE to capture the missing logical relations.
The key feature of logical bias terms is that they should be able to discriminate different types of tokens. Let the type of the $i$-th token be $t_{i}$ from a finite type set of size $t$. In our study, we consider six types of tokens: parenthesis[(,)], entity[s,e,f], relation[r], conjunction[$\land$], disjunction[$\lor$], and negation[!]. Then the bias vector $\beta_{ij}^{l}$ ($l$ is either $K$ or $V$) is indexed from an trainable embedding bank $\omega^{l}\in\mathbb{R}^{t\times t\times n\times d}$ (consider the first three indices).
$\beta_{ij}^{l}=\omega^{l}[t_{i},t_{j},|i-j|],\quad l\in\{K,V\}.$ (8)
LogiRPE may also be applicable to broader tasks where tokens with logical semantics can be labeled.
6.2 Free-variable Pooling
Most existing transformer-based approaches Bai et al. (2023b); Xu et al. (2023) utilize the last layer hidden states of the first token(s) as the final query encoding, akin to the use of the $[CLS]$ token in the original BERT paper Devlin et al. (2019). However, we contend that the reasoning capabilities inherent in self-attention mechanisms should effectively aggregate the query information into free variables that semantically represent the answer sets of queries. Consequently, we adopt Free-Variable Pooling, wherein the final query encoding is derived through max pooling across the last-layer hidden states of these free variables.
6.3 Result and Ablation
The experimental results of TEGA and Transformer baselines are presented in Table 7. In all three datasets, TEGA substantially improved the performance and generalizability of knowledge and query type dimensions over baselines. According to our ablation study results in Table 8, both inductive biases proposed in TEGA are effective on their own. At the same time, a stronger improvement in entailment performance can be achieved when applied together. Moreover, we provide an example of attention visualization in Appendix D, demonstrating the self-evident effect of logic-aware attention guidance within our TEGA model.
Dataset Model ID (Q) OOD (Q)
ID (K) OOD (K) ID (K) OOD (K)
FB15k Trans.+Absolute PE 46.9 31.9 21.8 13.2
Trans.+Relative PE 48.1 32.3 35.4 21.5
TEGA (ours) 55.0 34.1 38.8 22.4
FB15k-237 Trans.+Absolute PE 54.4 19.9 20.8 9.0
Trans.+Relative PE 54.8 20.0 37.7 15.8
TEGA (ours) 64.3 20.1 42.6 16.0
NELL995 Trans.+Absolute PE 68.8 18.9 25.8 8.3
Trans.+Relative PE 68.0 18.7 49.8 13.4
TEGA (ours) 75.0 19.2 56.6 13.7
Table 7: Experimental results of TEGA comparing to baselines in FB15k, FB15k-237 and NELL995.
Model ID (Q) OOD (Q)
ID (K) OOD (K) ID (K) OOD (K)
TEGA (ours) 55.0 34.1 38.8 22.4
[…]
Trans.+Relative PE (baseline) 48.1 32.3 35.4 21.5
Table 8: Ablation study on two inductive biases introduced in TEGA on FB15k.
7 Related Work
Transformers in Logical Reasoning
Transformers have exhibited remarkable performance in various forms of logical reasoning. In natural language inference (NLI) tasks Bowman et al. (2015); Williams et al. (2018), transformers analyze the logical relationship between a premise and a hypothesis, employing deductive reasoning over statements. Furthermore, transformers have been employed for inductive reasoning within rule-based systems Clark et al. (2020). Additionally, transformers have been applied to perform abductive reasoning in both natural language Bhagavatula et al. (2020) and formal language Bai et al. (2024); Gao et al. (2025) settings. Meanwhile, various techniques have been proposed to enhance logical reasoning in language understanding Chen (2023); Pan et al. (2023). Recent studies show that large language models (LLMs) exhibit enhanced logical reasoning abilities Parmar et al. (2024); Li et al. (2025); Fan et al. (2025), even without intermediate reasoning steps Zheng et al. (2025a). Logical entailment also plays a key role in hypothesis verification of LLMs, in areas such as commonsense reasoning Zhao et al. (2024), analogical reasoning Zheng et al. (2025b), and even scientific discovery Takagi et al. (2023); Zheng et al. (2025c).
KGQA with Parameterized Knowledge
Prior to this work, extensive research has been conducted on KGQA with parameterized knowledge. Regarding modeling approaches, iterative neural query encoders Ren and Leskovec (2020); Chen et al. (2022); Tsang et al. (2025) design representations for entity sets and execute logical operators iteratively, following the computational graph. In addition, neural-symbolic methods Zhu et al. (2022); Bai et al. (2023c); Yin et al. (2023b) incorporate link predictors and search over the symbolic space. In terms of the knowledge domain, existing benchmarks Ren et al. (2020); Ren and Leskovec (2020); Yin et al. (2023a) primarily utilize knowledge graphs containing general factual knowledge, while some studies have focused on commonsense knowledge Fang et al. (2024) and eventuality knowledge Bai et al. (2023a). Such queries can be extended to natural language settings with template or LLM-based approaches Zheng et al. (2024, 2025d); Zong et al. (2025); Bai et al. (2025a). The setting of KGQA with parameterized knowledge has also been extended to the concept of Neural Graph Databases Ren et al. (2023); Bai et al. (2025b).
8 Conclusion
This paper investigates the generalizable first-order logical reasoning capabilities of transformers when knowledge is parameterized within model weights. By establishing connections between distribution shifts and logical entailment, we develop a principled framework for evaluating transformer performance on knowledge graph query answering tasks. Our comprehensive benchmark reveals that transformers can effectively perform first-order logical entailment, outperforming methods specifically designed for this task. Through systematic analysis of the entire modeling pipeline, we identify critical design principles that enhance reasoning capabilities. Notably, our findings reveal a mismatch between existing inductive biases and optimal positional encoding strategies, motivating the development of TEGA. This architecture introduces logic-aware self-attention mechanisms that substantially improve both performance and generalizability. Our work demonstrates that careful architectural choices, informed by the logical structure of reasoning tasks, can significantly enhance transformers’ logical entailment ability.
Limitations
The scope of complex logical queries was limited to EFO queries with a single free variable, with queries containing multiple free variables reserved for future investigation. While LogiRPE incorporates fine-grained logic awareness, it remains dependent on explicit labeling. Future research should explore more generalizable approaches to logic-aware labeling that can accommodate previously unseen logical operators. With respect to knowledge graph scale, the largest knowledge graph employed in our experiments was FB15k, which is considered medium-sized within the knowledge graph literature. Consequently, the scalability of the proposed approaches to large-scale knowledge graphs remains an open question for future investigation.
Ethics Statement
The experiments were conducted on publicly available knowledge graphs, eliminating any data privacy concerns. However, it should be noticed that most approaches for generalizable logical entailment are susceptible to adversarial attacks Dai et al. (2018); Zügner et al. (2019) and data poisoning You et al. (2023) on knowledge graphs, which may result in unintended outcomes in applications.
Acknowledgements
We thank all the anonymous reviewers and meta reviewers for their valuable comments. The authors of this paper were supported by the ITSP Platform Research Project (ITS/189/23FP) from ITC of Hong Kong, SAR, China, and the AoE (AoE/E-601/24-N), the RIF (R6021-20) and the GRF (16205322) from RGC of Hong Kong, SAR, China.
Appendix A Dataset Statistics
In this section, we introduce the knowledge graph details and the statistics and query types of our benchmark dataset.
Our benchmark incorporates three widely used knowledge graphs—FB15k, FB15k-237, and NELL995—derived from large-scale KGs to evaluate reasoning capabilities. FB15k and FB15k-237 originate from Freebase, a comprehensive database of structured general knowledge covering domains such as entertainment, sports, and geography, with entities linked by diverse relations (e.g., "born-in," "part-of"). NELL995 is derived from the Never-Ending Language Learning (NELL) system, which extracts structured facts from web text, focusing on real-world entities and relations.
Details of the query statistics for our dataset are presented in Table 9. For each knowledge graph, we sampled all one-hop projection (1p) queries from the training graph and sampled twice as many for the other 22 in-distribution query types. For validation and testing queries, we set the quantities to 8000, 5000, and 4000, respectively, following the convention established in KG literature Ren et al. (2020); Ren and Leskovec (2020); Bai et al. (2023b).
Knowledge Graph Training Validation Testing
1p Other Types All Types All Types
FB15k 273,710 547,420 8,000 8,000
FB15k-237 149,689 299,378 5,000 5,000
NELL995 107,982 215,964 4,000 4,000
Table 9: Number of queries used for each query type in our benchmark.
Table 11 and 12 presented the details of query types in our benchmark. All 32 unseen query types are crafted compositionally based on the 23 seen query types. Table 13 presented the subset of 25 lisp-like compatible query types used in our experiment on query syntax.
Appendix B Baseline Models
In our benchmarking experiment, we compare transformers with five existing methods that are designed particularly for KG query answering:
  • BetaE Ren and Leskovec (2020) encodes queries iteratively following the logical operations with Beta distributions.
  • ConE Zhang et al. (2021) represents answer sets as two-dimensional cones to better handle negations.
  • CQD Arakelyan et al. (2021) applies a search-based method that utilizes pre-trained link predictors for inference-time optimization.
  • LMPNN Wang et al. (2023b) conducts one-hop inferences with pre-trained KGE, and uses message passing to aggregate these local results into global output.
  • SQE-LSTM Bai et al. (2023b) encode the linearized complex query as a sequence with LSTM Hochreiter and Schmidhuber (1997).
Appendix C Query Reversion
The logical formula of five query types before and after reversion is provided in Table 10.
Type Name Before Reversion After Reversion
2p (r1(s1,e1))&(r2(e1,f)) (r2(e1,f))&(r1(s1,e1))
3p ((r1(s1,e1))&(r2(e1,e2)))&(r3(e2,f)) ((r3(e2,f))&(r2(e1,e2)))&(r1(s1,e1))
ip ((r1(s1,e1))&(r2(s2,e1)))&(r3(e1,f)) ((r3(e1,f))&(r2(s2,e1)))&(r1(s1,e1))
pi ((r1(s1,e1))&(r2(e1,f)))&(r3(s2,f)) ((r3(s2,f))&(r2(e1,f)))&(r1(s1,e1))
2in (r1(s1,f))&(!(r2(s2,f))) (!(r2(s2,f)))&(r1(s1,f))
Table 10: Details of query reversion procedure in the query permutation experiment.
Appendix D Attention Visualization
Fig. 3 provides an example of how attention guidance in TEGA can affect the reasoning process in self-attention. In 2in queries [(r1(s1,f))&(!(r2(s2,f)))], the negation operator [!] is applied to calculate the complement of the entity set represented by the second atomic formula [(r2(s2,f))]. Benefited from the attention guidance from inductive biases, TEGA can effectively aggregate the information from the second atomic formula [r2/s2/f] to the negation operator with self-attention. While in baseline (Transformer + Relative PE), the information flow between tokens are scattered and less effective (highlighted in red). As a result, TEGA outperforms Transformer + Relative PE by 21-26% in 2in queries over three knowledge graphs.
Figure 3: An visualization of self-attention when processing 2in queries. The left and right image is the partial attention weight of TEGA (ours) and Transformer + Relative PE (baseline), respectively.
id type name mul. cyc. exi. neg. depth logical formula
0 1p 1 r1(s1,f)
1 2p 2 (r1(s1,e1))&(r2(e1,f))
2 3p 3 (r1(s1,e1))&(r2(e1,e2))&(r3(e2,f))
3 2i 1 (r1(s1,f))&(r2(s2,f))
4 3i 1 (r1(s1,f))&(r2(s2,f))&(r3(s3,f))
5 ip 2 (r1(s1,e1))&(r2(s2,e1))&(r3(e1,f))
6 pi 2 (r1(s1,e1))&(r2(e1,f))&(r3(s2,f))
7 2in 1 (r1(s1,f))&(!(r2(s2,f)))
8 3in 1 (r1(s1,f))&(r2(s2,f))&(!(r3(s3,f)))
9 inp 2 (r1(s1,e1))&(!(r2(s2,e1)))&(r3(e1,f))
10 pin 2 (r1(s1,e1))&(r2(e1,f))&(!(r3(s2,f)))
11 pni 2 (r1(s1,e1))&(!(r2(e1,f)))&(r3(s2,f))
12 2u 1 (r1(s1,f))|(r2(s2,f))
13 up 2 ((r1(s1,e1))|(r2(s2,e1)))&(r3(e1,f))
14 2m 2 (r1(s1,e1))&(r2(e1,f))&(r3(e1,f))
15 2nm 2 (r1(s1,e1))&(r2(e1,f))&(!(r3(e1,f)))
16 3mp 3 (r1(s1,e1))&(r2(e1,e2))&(r3(e2,f))&(r4(e1,e2))
17 3pm 3 (r1(s1,e1))&(r2(e1,e2))&(r3(e2,f))&(r4(e2,f))
18 im 2 (r1(s1,e1))&(r2(s2,e1))&(r3(e1,f))&(r4(e1,f))
19 2il 1 (r1(s1,f))&(r2(e1,f))
20 3il 1 (r1(s1,f))&(r2(s2,f))&(r3(e1,f))
21 3c 3 (r1(s1,e1))&(r2(e1,f))&(r3(s2,e2))&(r4(e2,f))&(r5(e1,e2))
22 3cm 3 (r1(s1,e1))&(r2(e1,f))&(r3(s2,e2))&(r4(e2,f))&(r5(e1,e2))&(r6(e1,f))
Table 11: Details of 23 seen query types with their type name, query-level features: mul:multiple relation projection edges, cyc:cycles, exi:existentially quantified variables, neg:negation, depth:longest relation projection chain (reasoning depth), and their logical formula.
id type name mul. cyc. exi. neg. depth logical formula
23 2pi 2 (r1(s1,e1))&(r2(e1,f))&(r3(s2,e2))&(r4(e2,f))
24 2pu 2 ((r1(s1,e1))&(r3(e1,f)))|((r2(s2,e2))&(r4(e2,f)))
25 ui 2 ((r1(s1,e1))|(r2(s2,e1)))&(r3(e1,f))&(r4(s3,f))
26 iu 2 ((r1(s1,e1))&(r2(s2,e1))&(r3(e1,f)))|(r4(s3,f))
27 upi 2 ((r1(s1,e1))|(r2(s2,e1)))&(r3(e1,f))&(r4(s3,e2))&(r5(e2,f))
28 ipu 2 ((r1(s1,e1))&(r2(s2,e1))&(r3(e1,f)))|((r4(s3,e2))&(r5(e2,f)))
29 i2p 3 (r1(s1,e1))&(r2(s2,e1))&(r3(e1,e2))&(r4(e2,f))
30 u2p 3 ((r1(s1,e1))|(r2(s2,e1)))&(r3(e1,e2))&(r4(e2,f))
31 2pin 2 (r1(s1,e1))&(r2(e1,f))&(r3(s2,e2))&(r4(e2,f))&(!(r5(s3,f)))
32 2pni 2 (r1(s1,e1))&(r2(e1,f))&(r3(s2,e2))&(!(r4(e2,f)))
33 pn3i 2 (r1(s1,e1))&(!(r2(e1,f)))&(r3(s2,f))&(r4(s3,f))
34 in2p 3 (r1(s1,e1))&(!(r2(s2,e1)))&(r3(e1,e2))&(r4(e2,f))
35 inu 2 ((r1(s1,e1))&(!(r2(s2,e1)))&(r3(e1,f)))|(r4(s3,f))
36 inpu 2 ((r1(s1,e1))&(!(r2(s2,e1)))&(r3(e1,f)))|((r4(s3,e2))&(r5(e2,f)))
37 upni 2 ((r1(s1,e1))|(r2(s2,e1)))&(r3(e1,f))&(r4(s3,e2))&(!(r5(e2,f)))
38 unpi 2 ((r1(s1,e1))|(r2(s2,e1)))&(!(r3(e1,f)))&(r4(s3,e2))&(r5(e2,f))
39 imp 3 (r1(s1,e1))&(r2(s2,e1))&(r3(e1,e2))&(r4(e2,f))&(r5(e1,e2))
40 ipm 3 (r1(s1,e1))&(r2(s2,e1))&(r3(e1,e2))&(r4(e2,f))&(r5(e2,f))
41 3im 2 (r1(s1,e1))&(r2(s2,e1))&(r3(s3,e1))&(r4(e1,f))&(r5(e1,f))
42 pil 2 (r1(s1,e1))&(r2(e1,f))&(r3(e2,f))
43 ilp 2 (r1(s1,e1))&(r2(e2,e1))&(r3(e1,f))
44 p3il 2 (r1(s1,e1))&(r2(e1,f))&(r3(s2,e2))&(r4(e2,f))&(r5(e3,f))
45 i3c 3 (r1(s1,e1))&(r2(s2,e1))&(r3(e1,f))&(r4(s3,e2))&(r5(e2,f))&(r6(e1,e2))
46 i3cm 3 (r1(s1,e1))&(r2(s2,e1))&(r3(e1,f))&(r4(s3,e2))&(r5(e2,f))&(r6(e1,e2))&(r7(e1,f))
47 3inl 1 (r1(s1,f))&(!(r2(s2,f)))&(r3(e1,f))
48 pinl 2 (r1(s1,e1))&(r2(e1,f))&(!(r3(s2,f)))&(r4(e2,f))
49 inm 2 (r1(s1,e1))&(!(r2(s2,e1)))&(r3(e1,f))&(r4(e1,f))
50 inmp 3 (r1(s1,e1))&(!(r2(s2,e1)))&(r3(e1,e2))&(r4(e2,f))&(r5(e1,e2))
51 inpm 3 (r1(s1,e1))&(!(r2(s2,e1)))&(r3(e1,f))&(r4(e2,f))&(r5(e2,f))
52 3nmp 3 (r1(s1,e1))&(r2(e1,e2))&(r3(e2,f))&(!(r4(e1,e2)))
53 3cn 3 (r1(s1,e1))&(r2(e1,f))&(r3(s2,e2))&(!(r4(e2,f)))&(r5(e1,e2))
54 3cnm 3 (r1(s1,e1))&(r2(e1,f))&(r3(s2,e2))&(r4(e2,f))&(r5(e1,e2))&(!(r6(e1,f)))
Table 12: Details of 32 unseen query types with their type name, query-level features: mul:multiple relation projection edges, cyc:cycles, exi:existentially quantified variables, neg:negation, depth:longest relation projection chain (reasoning depth), and their logical formula.
Seen Query Types (13 types)
Type Name EFO Syntax Lisp-Like Syntax
1p r1(s1,f) (p,(r1),(s1))
2p (r1(s1,e1))&(r2(e1,f)) (p,(r2),(p,(r1),(s1)))
3p (r1(s1,e1))&(r2(e1,e2))&(r3(e2,f)) (p,(r3),(p,(r2),(p,(r1),(s1))))
2i (r1(s1,f))&(r2(s2,f)) (i,(p,(r1),(s1)),(p,(r2),(s2)))
3i (r1(s1,f))&(r2(s2,f))&(r3(s3,f)) (i,(p,(r1),(s1)),(p,(r2),(s2)),(p,(r3),(s3)))
ip (r1(s1,e1))&(r2(s2,e1))&(r3(e1,f)) (p,(r3),(i,(p,(r1),(s1)),(p,(r2),(s2))))
pi (r1(s1,e1))&(r2(e1,f))&(r3(s2,f)) (i,(p,(r2),(p,(r1),(s1))),(p,(r3),(s2)))
2in (r1(s1,f))&(!(r2(s2,f))) (i,(p,(r1),(s1)),(n,(p,(r2),(s2))))
3in (r1(s1,f))&(r2(s2,f))&(!(r3(s3,f))) (i,(p,(r1),(s1)),(p,(r2),(s2)),(n,(p,(r3),(s3))))
inp (r1(s1,e1))&(!(r2(s2,e1)))&(r3(e1,f)) (p,(r3),(i,(p,(r1),(s1)),(n,(p,(r2),(s2)))))
pin (r1(s1,e1))&(r2(e1,f))&(!(r3(s2,f))) (i,(p,(r2),(p,(r1),(s1))),(n,(p,(r3),(s2))))
2u (r1(s1,f))|(r2(s2,f)) (u,(p,(r1),(s1)),(p,(r2),(s2)))
up ((r1(s1,e1))|(r2(s2,e1)))&(r3(e1,f)) (p,(r3),(u,(p,(r1),(s1)),(p,(r2),(s2))))
Unseen Query Types (12 types)
Type Name EFO Syntax Lisp-Like Syntax
2pi (r1(s1,e1))&(r2(e1,f))&(r3(s2,e2))&(r4(e2,f)) (i,(p,(r2),(p,(r1),(s1))),(p,(r4),(p,(r3),(s2))))
2pu ((r1(s1,e1))&(r3(e1,f)))|((r2(s2,e2))&(r4(e2,f))) (u,(p,(r3),(p,(r1),(s1))),(p,(r4),(p,(r2),(s2))))
ui ((r1(s1,e1))|(r2(s2,e1)))&(r3(e1,f))&(r4(s3,f)) (i,(p,(r4),(s3)),(p,(r3),(u,(p,(r1),(s1)),(p,(r2),(s2)))))
iu ((r1(s1,e1))&(r2(s2,e1))&(r3(e1,f)))|(r4(s3,f)) (u,(p,(r4),(s3)),(p,(r3),(i,(p,(r1),(s1)),(p,(r2),(s2)))))
upi ((r1(s1,e1))|(r2(s2,e1)))&(r3(e1,f))&(r4(s3,e2))&(r5(e2,f)) (i,(p,(r5),(p,(r4),(s3))),(p,(r3),(u,(p,(r1),(s1)),(p,(r2),(s2)))))
ipu ((r1(s1,e1))&(r2(s2,e1))&(r3(e1,f)))|((r4(s3,e2))&(r5(e2,f))) (u,(p,(r5),(p,(r4),(s3))),(p,(r3),(i,(p,(r1),(s1)),(p,(r2),(s2)))))
i2p (r1(s1,e1))&(r2(s2,e1))&(r3(e1,e2))&(r4(e2,f)) (p,(r4),(p,(r3),(i,(p,(r1),(s1)),(p,(r2),(s2)))))
u2p ((r1(s1,e1))|(r2(s2,e1)))&(r3(e1,e2))&(r4(e2,f)) (p,(r4),(p,(r3),(u,(p,(r1),(s1)),(p,(r2),(s2)))))
2pin (r1(s1,e1))&(r2(e1,f))&(r3(s2,e2))&(r4(e2,f))&(!(r5(s3,f))) (i,(i,(p,(r4),(p,(r3),(s2))),(p,(r2),(p,(r1),(s1)))),(n,(p,(r5),(s3))))
in2p (r1(s1,e1))&(!(r2(s2,e1)))&(r3(e1,e2))&(r4(e2,f)) (p,(r4),(p,(r3),(i,(p,(r1),(s1)),(n,(p,(r2),(s2))))))
inu ((r1(s1,e1))&(!(r2(s2,e1)))&(r3(e1,f)))|(r4(s3,f)) (u,(p,(r4),(s3)),(p,(r3),(i,(p,(r1),(s1)),(n,(p,(r2),(s2))))))
inpu ((r1(s1,e1))&(!(r2(s2,e1)))&(r3(e1,f)))|((r4(s3,e2))&(r5(e2,f))) (u,(p,(r5),(p,(r4),(s3))),(p,(r3),(i,(p,(r1),(s1)),(n,(p,(r2),(s2))))))
Table 13: Details of 25 query types (compatible in both Lisp-like and EFO syntax) that we used as a subset for exploration on query syntax.
Appendix E Knowledge Graph Statistics
Detailed statistics of the knowledge graphs selected are presented in Table 14.
Dataset Relations Entities Training Validation Testing All Edges
FB15k 1,345 14,951 483,142 50,000 59,071 592,213
FB15k-237 237 14,505 272,115 17,526 20,438 310,079
NELL995 200 63,361 114,213 14,324 14,267 142,804
Table 14: Details of three knowledge graphs used for the experiments, and their separation standard for training, validation, and testing edges according to Ren and Leskovec (2020).
Appendix F Query Graph Definition
We provide query graph for each query types in our dataset. Query graphs can be utilized for graph-augmented methods, and have been applied in the two inductive biases discussed in Sec. 5.3. Here we provide the definition:
For each atomic formula or its negation $\alpha=r(h,t)$ or $\neg r(h,t)$ in a conjunctive formula $c$, we have $\{(h,r),(r,t)\}\in\mathcal{G}_{c}$ or $\{(h,r),(r,n),(n,t)\}\in\mathcal{G}_{c}$, where $n$ denotes the negation node in the conjunctive query graph $\mathcal{G}_{c}$. By our definition of query, all conjunctive query graphs have exactly one node as a free variable. For queries that represent the disjunction of multiple conjunctive formulas, we replace all free variable nodes in every conjunctive query graph with a single union node $u$, and connect it to a final free variable node $f$, with $\{(u,f)\}\in\mathcal{G}_{d}$, where $\mathcal{G}_{d}$ is the disjunctive query graph.
Appendix G Evaluation Details
For each knowledge graphs, we separate their edges into training, validation, and testing edges with a ratio of approximately 8:1:1, as shown in Table 14. We construct three graphs, training graph $\mathcal{G}_{train}$, validation graph $\mathcal{G}_{valid}$, and testing graph $\mathcal{G}_{test}$ with training edges, training+validation edges, and training+validation+testing edges, respectively.
We adopt the mean reciprocal rank (MRR) as our evaluation metric, following the calculation presented in Equation 9 and 10. It is important to note that for all testing queries, the answers in $A(\mathcal{G}_{valid},q)-A(\mathcal{G}_{train},q)$ are excluded from $\mathcal{A}_{ood}$ to ensure fairness. Consequently, we define $\mathcal{A}_{id}=A(\mathcal{G}_{train},q)$ and $\mathcal{A}_{ood}=A(\mathcal{G}_{test},q)-A(\mathcal{G}_{valid},q)$ for our evaluation of ID (K) and OOD (K), respectively. In each table regarding experimental results, the scores below ID (Q) (i.e., either MRR of ID (K) or OOD (K)) represent the average scores among all seen query types. While the scores below OOD (Q) represent the average scores among all unseen query types. To connect with literature in KG query answering Bai et al. (2023b); Sun et al. (2021), the performances in ID (K) and OOD (K) are also referred to as faithfulness and knowledge inference capability, respectively.
$MRR_{id(k)}$ $=\frac{\sum_{v\in\mathcal{A}_{id}}}{|\mathcal{A}_{id}|\texttt{rank}(v)}$ (9)
$MRR_{ood(k)}$ $=\frac{\sum_{v\in\mathcal{A}_{ood}}}{|\mathcal{A}_{ood}|\texttt{rank}(v)}$ (10)
Appendix H Transformer Architecture Comparison
We investigate the impact of different transformer architectures on complex query answering performance. Table 15 compares three architectural variants: encoder-only (our main approach), decoder-only, and encoder-decoder configurations. The results demonstrate that the encoder-only architecture with relative positional encoding achieves the best overall performance across datasets.
Dataset Architecture PE Type ID (Q) OOD (Q) All Queries
ID (K) OOD (K) ID (K) OOD (K) ID (K) OOD (K)
[…]
Encoder-only Relative PE 48.1 32.3 35.4 21.5 40.7 26.0
[…]
FB15k-237 Encoder-only Absolute PE 54.4 20.0 20.8 9.0 34.9 13.6
[…]
Decoder-only Absolute PE 55.2 19.9 21.9 9.8 35.8 14.0
[…]
Decoder-only Absolute PE 65.6 18.6 28.0 8.9 43.7 13.0
[…]
Table 15: Comparison of different transformer architectures (in MRR%).
Appendix I Technical Details
In this section, we describe our experiment setting in more details.
Computational Resource: All transformers are trained on four NVIDIA A100 GPUs for two days, with a batch size of 1024. As shown in Table 16, TEGA achieves computational efficiency comparable to that of the baseline models.
Model Configuration: Table 18 illustrated the impact of transformer size to query answering performances. We selected a three-layer configuration for all Transformer models, for its effectiveness and comparable parameter size to other baseline methods (as shown in Table 17). To train the Transformer models, we employed the label smoothing loss Szegedy et al. (2015) with a smoothing factor of $\epsilon=0.1$, which helps to regularize the model and improve generalization. The learning rate was set to 0.0001, and a warm-up schedule of 1000 steps was applied to gradually increase the learning rate during the initial phase of training, allowing the model to adapt to the task and stabilize the optimization process. All Transformer models and baselines were configured with an embedding size of 400 dimensions. Regarding transformer encoder with absolute and relative PE, we employ the BertModel from the transformers library. While for disentangled PE, we use DebertaModel with the same size and configurations. Note that apart from disentangled PE, Deberta also proposed a mask decoder during the pre-training stage. However, as we train all models solely on logical queries without utilizing any pre-trained model weights, the impact of the mask decoder is negligible.
Pooling Strategy: For free-variable pooling, we experimented with two pooling settings: sum pooling and mean pooling, with results presented in Table 19. The results show that sum pooling over free variables is the better setting, which is also consistent with findings in GNN literature Xu et al. (2019).
Models Inference Time (ms)
Transformer (Absolute PE) 43.2
Transformer (Disentangled PE) 47.3
Transformer (Relative PE) 44.7
+Adjacency Matrices Masking 46.8
+Direct Distance Encoding 44.9
TEGA (ours) 52.1
Table 16: Computational efficiency of TEGA compared with baselines.
Model Number of Trainable Parameters (in Millions)
BetaE 13.7
ConE 18.9
LMPNN 11.5
SQE-LSTM 8.8
Trans.+Relative PE 14.7
TEGA (ours) 15.7
Table 17: Parameter size comparison between different query answering approaches, with three layers of self-attention in transformer models.
Model # Attn. layers ID(Q) OOD(Q) All Queries
ID(K) OOD(K) ID(K) OOD(K) ID(K) OOD(K)
[…]
6 57.2 20.3 23.0 9.6 37.3 14.1
12 48.9 19.8 24.4 11.4 34.6 14.9
[…]
12 48.6 19.9 35.0 15.7 40.7 17.4
Table 18: Experimental results on transformers with different numbers of attention layers in the FB15K-237 dataset.
Dataset Pooling ID (Q) OOD (Q) All Queries
ID (K) OOD (K) ID (K) OOD (K) ID (K) OOD (K)
FB15k Mean 52.9 33.5 37.2 21.6 43.7 26.6
Sum 55.0 34.1 38.8 22.4 45.6 27.3
FB15k-237 Mean 62.9 20.0 41.6 15.6 50.5 17.4
Sum 64.3 20.1 42.6 16.0 51.7 17.7
Table 19: Performance comparison between sum pooling and mean pooling strategies on TEGA.